Concord Inc. manufactures cycling equipment. Recently, the vice president of operations of the company has requested construction of a new plant to meet the increasing demand for the company’s bikes. After a careful evaluation of the request, the board of directors has decided to raise funds for the new plant by issuing $3,294,900 of 12% term corporate bonds on March 1, 2020, due on March 1, 2035, with interest payable each March 1 and September 1, with the first interest payment on September 1st, 2020. At the time of issuance, the market interest rate for similar financial instruments is 10%. Click here to view factor tables As the controller of the company, determine the selling price of the bonds.
Answer #1
Issue price of the bond=Present value of interest payment+Present value of bond at the time of maturity
Interest payment per semi-annual period=Face value of the bond*stated interest rate*1/2=3294900*12%*1/2=$ 197694
Discount factor=market rate for 6 months=10%*6/12=5%
Life of the bond=Mar 1,2020 to Mar 1,2035=15 years
Number of semi-annual periods=15*2=30
Present value of interest payment=197694*Present value at 5% for 30 years=197694*15.37245=$ 3039041
Present value of bond at the time of maturity=3294900*present value at 5% for 30th year=3294900*0.231377=$ 762364
Issue price of the bond=3039041+762364=$ 3801405
I appreciate your ratings